OpenStudy (anonymous):

A restaurant charges $11 for a buffet plus $1 for each drink. Another restaurant charges $9 for a buffet plus $2 for each drink. The system of equations is graphed below.

OpenStudy (anonymous):

Part 1: What is the solution? Part 2: Show your work to verify the solution for both equations.

OpenStudy (anonymous):

I'm not sure on this one.

OpenStudy (anonymous):

Me either, thanks though! ><

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

I figured it out ! ^^

OpenStudy (anonymous):

Great and thx for the medal!

OpenStudy (anonymous):

So what was the answer

OpenStudy (anonymous):

well you have to make two linear equations out of the information that you were given. then with those equations you have to find a number that works with both equations. 11 + 1d & 9 + 2d ----------------- 11 + 1(1) = 12 9 + 2(1) = 11 , no ----------------- 11 + 1(2) = 13 9 + 2(2) = 13 ----------------- So if two drinks were ordered at the first restaurant; 11 + 1(2) = 13 If two drinks were ordered at the first restaurant; 9 + 2(2) = 13 So the solution is 2.

OpenStudy (anonymous):

i guess in this case it would be two expressions xD

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

My dad is a calculus So he helps me alot